OpenStudy (anonymous):

solve the following: -3z+8=-3+4z

OpenStudy (anonymous):

This is very similar to the others. Do you want to try it first?

OpenStudy (anonymous):

I'm trying it.

OpenStudy (anonymous):

subract 8

OpenStudy (anonymous):

From both sides. Yes.

OpenStudy (anonymous):

-3z=-11+4z

OpenStudy (uri):

-3z+8=-3+4z -8 -8 =========== -3z=-11+4z

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

do I add 3z or subtract 4z????????????

HanAkoSolo (jamierox4ev3r):

-3z+8=-3+4z +3z+3 +3+3z ------------- 11=7z /7 /7 z=\(\frac{11}{7}\) This is just a mapout of the process, to help you understand

OpenStudy (anonymous):

You want the variable to be on the left so -4z from both sides because it is on the right. (Not where you want it).

OpenStudy (anonymous):

11/7??

OpenStudy (anonymous):

Yes. You may need to change the improper fraction to a mixed number depending on your teacher's requirements.

OpenStudy (uri):

-3z=-11+4z -4z-3z=-11 -7z=-11

OpenStudy (uri):

z=-11/-7 z=11/7